Main INDEX
Monthly INDEX
PREV
NEXT
Make New Entry,
Make Followup Entry
User name V. Sulkosky
Log entry time 17:42:25 on February 21, 2009
Entry number 262525
This entry is a followup to: 261625
Followups:
keyword=Experiment Summary for Feb 20, Owl - Feb 21, Owl
Summary for Feb 20 owl - Feb 21 owl
Smooth data taking.
Beam in Use
shift |
ABU |
BANU |
BNA |
2009/02/20 owl | 7.27 |
0.28 | 0.45 |
2009/02/20 day | 5.52 |
1.69 | 0.8 |
2009/02/20 swing | 6.33 |
1.00 | 0.68 |
2009/02/21 owl | 6.79 |
0.29 | 0.94 |
total | 25.9 |
3.2 | 2.9 |
Main events/issues:
- Transverse Production
- Feb. 20, 08:50 Completed kinematic J: Plhrs = 1.42
GeV/c, 45o.
- Feb. 20, 08:55 Start kinematic H: Plhrs = 1.42 GeV/c,
45o.
- Feb. 20, 13:30-14:30 Controlled access to fix network slow down.
- Feb. 20, 15:00 Resume production on kinematic H.
- Feb. 20, 19:05 Change target spin direction to 90 degrees.
- Average target polarization: ~ 62.5% from owl shift 2/20 through
owl shift 2/21.
- Total accummulated charge for kinematic J: 1.24 C out of 1.30 C.
- Total accummulated charge for kinematic H: 0.77 C out of 1.08 C.
- Issue: BigBite DAQ sees deadtime burps where the DAQ stops for
about 1-2 seconds.
Plans for the next day:
- Complete kinematic H: Plhrs = 1.42 GeV/c.
- Begin kinematic F: Plhrs = 1.13 GeV/c.
A copy of this log entry has been emailed to: brads,jpchen,jiang,meziani,choi